OpenStudy (anonymous):

find dy/dx when 3(square root of x) +2(square root of y)=5 What I did was (1/2)(3)x^(-1/2) + y^(-1/2)*dy/dx=0 After solving for dy/dx and simplifying I got -3(square root of y)/2(square root of x) However, the answers do not match up to this. Can someone help me? 1. dy/dx=(3/4)(5+3/(square root of x)) 2.dy/dx=5-3/(square root of x) 3.dy/dx=(3/4)(3-5/(square root of x)) 4.dy/dx=(3/2)(3-5/(square root of x)) 5. dy/dx=(3/2)(5-3/(square root of x)) 6.dy/dx=3+5/(square root of x))

OpenStudy (anonymous):

\[ 3\sqrt{x}+2\sqrt{y}=5 \]I'd first solve for \(y\) here.

OpenStudy (anonymous):

Or solve for just \(\sqrt{y}\)

OpenStudy (anonymous):

Notice how their answers only have \(x\) in them.

OpenStudy (anonymous):

Then you plug in your \(\sqrt{y}\) result into your implicit differentiation equation.